The velocity of ball of mass 100 gram changes from 20 m/s in 2 second. Find change in momentum and force acting on ball

Grade:9

1 Answers

Ayush Singh
5 years ago
Given mass=100gm=0.1kgVelocity 1=20m/secTime=2secFind-1-change in momentum2 forceChange in momentum=m(v2-v1)=0.1(x-20)=(0.1x-2)kgm/secForce=mass*acceltation=0.1*(x-20)/2=0.05x-1 Newton